Consider the following two projects:

Year                            Cash Flow (Beta)        Cash Flow (Zeta)

0                                  −$25,000                     −$28,000

1                                         12,000                  14,000

2                                         10,000                  13,000

3                                         9,000                    11,000

Instructions:

1. Using company cost of capital 15%, calculate the following investment criteria for both projects:

a.     Payback period                                               

b.     Net Present Value (NPV)                               

c.     Internal Rate of Return (IRR)                         

d.     Profitability Index (PI)                                     

2. If projects Beta and Zeta are independent, which one(s) will you choose? Why?            

3. If projects Beta and Zeta are mutually exclusive, which one will you choose? Why?
Answer #1

Using financial calculator to solve all parts of question 1

Q1a) Payback of zeta

Inputs: C0 = -28,000

C1 = 14,000. Frequency= 1

C2 = 13,000. Frequency= 1

C3= 11,000. Frequency= 1

Now press, Npv button and scroll down till PB, then press compute

We get, payback period of zeta as 2.09 years

Payback period for Beta

Inputs: C0 = -25,000

C1 = 12,000. Frequency= 1

C2 = 10,000. Frequency= 1

C3 = 9,000. Frequency= 1

Use the same step as above 2.33 years

We get, payback period of Beta as

B) NPV of Zeta

Inputs: C0 = -28,000

C1 = 14,000. Frequency= 1

C2 = 13,000. Frequency= 1

C3 = 11,000. Frequency= 1

I = 15%

Npv = compute

We get, NPV of zeta as $1,236.459

NPV of Beta

Inputs: C0 = -25,000

C1 = 12,000. Frequency= 1

C2 = 10,000. Frequency= 1

C3 = 9,000. Frequency= 1

I = 15%

Npv = compute

We get, NPV of Beta as -$1,086.13

C) IRR of Zeta

Inputs: C0 = -28,000

C1 = 14,000. Frequency= 1

C2 = 13,000. Frequency= 1

C3 = 11,000. Frequency= 1

Irr = compute

We get, IRR of zeta as 17.76%

IRR of Beta

Inputs: C0 = -25,000

C1 = 12,000. Frequency= 1

C2 = 10,000. Frequency= 1

C3 = 9,000. Frequency= 1

Irr = compute

We get, Irr of Beta as 12.23%

D) PI of zeta

PI = Npv + Cash outflow / initial investment

= 1,236.459 + 28,000 / 28,000

= 29,236.459 / 28,000

= 1.044

PI of Beta

PI = Npv + Cash outflow / initial investment

= -1,086.13 + 25,000 / 25,000

= 23,913.87 / 25,000

= 0.957

Q2) If projects are independent choose Zeta as it has positive NPV.

Q3) IF projects are mutually exclusive then also choose Zeta because it has higher Npv than beta.